OpenStudy (anonymous):

Find the following quotient 16x^2 - 9y^2 / 4x + 3y

OpenStudy (anonymous):

4x-3y

OpenStudy (anonymous):

just use a^2-b^2 =(a+b)*(a-b) identity

OpenStudy (anonymous):

so i am not using FOIL

OpenStudy (anonymous):

what is FOIL?

OpenStudy (anonymous):

first outer inner last

hero (hero):

16x^2 - 9y^2 / 4x + 3y ((4x+3y)(4x-3y))/(4x+3y) = 4x-3y

hero (hero):

I thought FOIL was Aluminum

OpenStudy (anonymous):

ya two year olds do it like that hahaha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

well when i suck i guess i have to do it like that lol

OpenStudy (anonymous):

i mean seriously did u use FOIL to solve the problem bella?

hero (hero):

FOIL sucks

OpenStudy (anonymous):

yes maybe okay yes

OpenStudy (anonymous):

stop laughing at me okay am 2

hero (hero):

No, it really does suck in real life.

OpenStudy (anonymous):

wait ur 2 , ur joking right?

hero (hero):

If she's really 2, then she's a genius

OpenStudy (anonymous):

not am not 2 am joking

OpenStudy (anonymous):

awe shucks

OpenStudy (anonymous):

Nah guys i have to do it this way i can pass all my classes without a sweat but algebra is gonna be the death of me

OpenStudy (anonymous):

Just keep trying ,and u will find ease at the end of the tunnel

OpenStudy (anonymous):

can you check something else for me please

OpenStudy (anonymous):

ya sure bella

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

(y^2 + 4y -7)(y+3)

OpenStudy (anonymous):

find the product

OpenStudy (anonymous):

y^3 + 4y^2 -7y +3y^2 +12y -21 which when simplified gives y^3 +7y^2 + 5y -21

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

man i didn't see it like that i didn't think to remember that 5 and 2 is 7

hero (hero):

Stom, you failed to show your steps. I wanted to see if you could solve it without FOIL

hero (hero):

I don't think 5 and 2 was used to get 7

OpenStudy (anonymous):

hey whats going on guys 5 +2 is 7 whats wrong in it??

OpenStudy (anonymous):

i didn't see anything wrong in it

hero (hero):

When did you add 5 and 2 to get 7?

OpenStudy (anonymous):

if anything Hero be a Hero and show a girl where i went wrong

OpenStudy (anonymous):

4+3=7

OpenStudy (anonymous):

awe darn missed that one

OpenStudy (anonymous):

do u want me to show the steps bell

OpenStudy (anonymous):

4y^2 + 3y^2 you add the like terms

OpenStudy (anonymous):

yes please it would help

hero (hero):